OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. x + y - 3z = -8 2x + 2y + z = 12 3x + y - z = -2 A. x + y = -8 3x + y = -2 B. x + y = -8 5x + 3y = 14 C. 7x + 7y = 28 3x + y = -2 D. 7x + 7y = 28 5x + 3y = 10

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@texaschic101 help me please

OpenStudy (texaschic101):

x + y - 3z = -8 2x + 2y + z = 12 --->(3) ------------ x + y - 3z = -8 6x + 6y + 3z = 36 (result of multiplying by 3) -------------add 7x + 7y = 28 2x + 2y + z = 12 3x + y - z = -2 ---------------add 5x + 3y = 10 solution is D